### 26 Jan 2024 v1

@@ 3.1 Merger histories and satellite fraction
In this subsection, we examine the merger histories of the three descendant samples (disc, lenticular and spheroidal) in which mergers have been identified, as discussed in section 2.5.
We examine the cumulative distribution of major and minor mergers experienced by galaxies from redshift $z=1$ (corresponding to approximately 8 Gyrs ago) to redshift $z=0$. The analysis is conducted on the descendant galaxy samples as depicted in Fig 4. Overall, descendant galaxies with a disc morphology have relatively quiet merger histories. Specifically, $85$ per cent of these galaxies have not undergone a major merger in the past $8$ Gyrs. In contrast, only $38$ per cent and $36$ per cent of the descendant lenticular and spheroidal galaxies, respectively, are free of major mergers. This is in agreement with previous theoretical studies (e.g., see Fig.16 in Izquierdo-Villalba et al. 2019). Additionally, in the second panel of Fig. 4 we appreciate that disc descendant galaxies that have undergone a significant merger, on average, experience the last major merger event approximately $4.7$ Gyr ago, but lenticular and spheroidal descendant galaxies typically undergo later their last large merger at $3.7$ and $3.8$ Gyr ago, respectively.
Regarding the minor merger histories, no notable distinction is observed in the cumulative distribution across descendant galaxies of different types, with more than 30 per cent of them experiencing at least one minor merger. Specifically, approximately $62$ per cent of disc and lenticular and $66$ per cent of spheroidal descendant galaxies do not undergo minor mergers. The lenticulars underwent their most recent minor merger on average $3.9$ Gyrs ago. In contrast, the disc and lenticular descendant galaxies exhibited earlier occurrences of minor mergers, thereby confirming the notion of relatively quiet evolutionary paths for these types of galaxies. The data presented in Table 2 illustrates the fraction of galaxies that have undergone at least one significant merger event since redshift $z=1$ across various samples. Additionally, the table provides the median values for the most recent major and minor merger events within each sample.
Figure 4: The merger histories of the three galaxy samples as specified in the legend. From left to right panels: The cumulative distribution of major mergers, the cumulative distribution of the last major merger, the cumulative distribution of the minor mergers, and the cumulative distribution of the last minor merger. Overall, the descendant disc galaxies have more quiet major merger histories when compared to lenticular and spheroidal descendant galaxies.
Table 2: From left to right columns: fraction of galaxies with at least one major mergers since $z=1$, fraction of galaxies with at least one minor merger and Median lookback time of last major ($t_{\rm Mm}$[Gyr]), minor ($t_{\rm mm}$[Gyr]) merger.
@@
One significant element that may be playing a role in this morphological transformation is the influence of tidal torques generated by the dense environment in which these galaxies are located (e.g. Martinez-Valpuesta et al. 2016; Łokas et al. 2016; Łokas 2021; Méndez-Abreu, Costantin, & Kruk 2023). This influence becomes particularly pronounced when these galaxies transition into denser environments. One approach to gaining insight into this phenomenon is by an examination of the evolution of satellite fraction for each descendant sample, since satellite galaxies are susceptible to experiencing heightened levels of tidal torques, ram pressure stripping, and mergers.
@@
Fig. 5 displays the fraction of satellites in the parent sample and each descendant sample. Overall, the satellite fraction increases with time. The highest increase is presented by the spheroidal descendants, with satellite galaxies accounting for $20$ per cent at redshift $z=1$ and expanding to $50$ per cent in the current epoch. Disc and lenticular descendants also present an increase, albeit of a more modest kind. Specifically, the fraction of galaxies classified as satellites is at less than $10$ per cent at redshift $z=1$, gradually escalating to reach approximately 40 per cent. These galaxies are marked by a higher occurrence of interactions and a larger degree of exposure to tidal forces. Several processes in denser environments have the capacity to either disrupt or create a bar structure. This topic will be further examined in the following sections.
@@
Figure 5: The redshift evolution of the satellite fraction for the parent sample and the disc, lenticular and spheroidal descendant samples. The satellite fraction increases with decreasing redshift, with the highest evolution presented for the spheroidal descendants.
4 The evolution of the $z=1$ barred galaxies
@@
In this section, we will explore the evolution of the bar structures in the barred galaxies at $z=1$. For this, we will focus only on disc and lenticular galaxies. As discussed in the previous section, disc descendant galaxies, despite having generally quiet merger histories, exhibit a diversity of merger histories and environments (see Fig 4), which must be taken into account when investigating the evolution of bar structures. For the lenticular galaxies, this diversity in the merger histories is higher. Each barred galaxy in the disc and lenticular descendants has been divided into 2 subsamples: (1) the descendants of barred galaxies at $z=1$ remain as barred galaxies at $z=0$ ;(2) the descendants of barred galaxies at $z=1$ lose their bar, becoming unbarred galaxies at $z=0$.
@@
Figure 6: The evolution of barred disc galaxies divided into 2 subsamples: barred galaxies at $z=1$ which remain barred galaxies at $z=0$ (left column), and those that do not have their bar at $z=0$ (right column). Solid lines represent the median values and the shaded area corresponds to the $20^{\rm th}$ and $80^{\rm th}$ percentiles of the distribution.
@@
Figure 7: The evolution of lenticular descendant galaxies divided into 2 subsamples: barred galaxies at $z=1$ which remain barred galaxies at $z=0$ (left column), those that do not have their bar at $z=0$ (right column). The evolution of disc galaxies is shown in Fig. 6. Solid lines represent the median values and the shaded area corresponds to the $20^{\rm th}$ and $80^{\rm th}$ percentiles of the distribution. Note that here we only have one lenticular galaxy that loses its bar.
@@
In our analysis, which is summarised in Table 1, it is shown that a total of $6$ galaxies, or $3$ per cent of the parent sample (the $z=1$ disc galaxies), experience the loss of the bar structure. On the other hand, a third of the galaxies in our parent sample, namely $56$ barred galaxies or $30$ per cent (including bars only in lenticular and disc galaxies) keep a bar structure at $z=0$. Upon examination of each descendant barred galaxy sample, a consistent pattern emerges: A minority of disc and lenticular descendent galaxies experienced bar dissolution, with fractions of $4.7$ per cent and $1.1$ per cent respectively. Conversely, the majority of these galaxies maintained their existing bar.
@@
The evolution of the properties of the bar structures is illustrated in Fig. 6 and Fig. 7 over the barred disc and lenticular galaxies, respectively. The first columns of the figures depict the evolution of the bar structures in the descendant galaxies that have retained their bars until the present epoch ($z=0$). We can appreciate that the bars that were already formed at $z=1$, formed at an early time (approximately $10$ Gyrs ago), and the evolution of their bars follows a typical evolution of a bar observed in isolated galaxies. Previous studies (Martinez-Valpuesta, Shlosman, & Heller 2006), have shown that bars experience rapid growth during their dynamical phase that lasts the first Gyrs, after which their evolution becomes secular and remains relatively constant. However, there is a possibility that the bars may slightly weaken over time due to a buckling instability, as explained by Martinez-Valpuesta, Shlosman, & Heller (2006) and then remain constant in its secular evolution. This instability could potentially account for some of the fluctuations observed in the length of bars. Upon examining the morphological evolution depicted in the third and fourth rows of Fig. 6 & Fig. 7, it becomes evident that there is a lack of significant evolution in the thick disc fraction. However, a marginal drop in the thin disc fraction is observed (also seen in the evolution of the lenticular sample). This phenomenon is associated with the rise in the bulge mass fraction ($B/T$) of both classical bulges and pseudobulges during later times (see also Izquierdo-Villalba et al. 2022). This increase could likely be attributed to the presence of a bar, which facilitates the inflow of gas toward the central regions of galaxies (spinoso2017; donohoe2019; George et al. 2019; Fraser-McKelvie et al. 2020) or as a result of mergers, although, in the next section, we will see that barred galaxies of the disc have the quietest merger histories. This increase in the $B/T$ is also found in lenticular galaxies, and the magnitude of this increase is greater. It is important to observe that the bars from the descendant barred disc galaxies that have retained their bar structure are more elongated than those in lenticular galaxies.
@@
The evolution of the bar structures that are dissolved is shown in the right columns of Fig. 6 and Fig. 7. These galaxies under consideration constitute a smaller fraction, specifically $4.7$ and $1.1$ per cent, of the total population of disc and lenticular descendants, respectively. It should be noted that the formation of bars occurred just a prior time steps of $z=1$. These bars exhibit relatively lower strength and shorter length when compared to those in galaxies that conserved their bar. Regarding morphology, the disc mass fraction remains rather stable at $z=0$ for the disc galaxy samples, with occasional changes likely attributed to either a major or minor merger event. The bulge mass fraction exhibits a noticeable rise but with a significant degree of scatter. Additionally, it is important to note that the thick disc and pseudobulge components do not exhibit any signs of evolutionary changes. In the case of the lenticular galaxies that lose their bar (only one case), not only the bar structures but the thin disc is disrupted whereas there is not an increase in the classical and Pseudo bulge mass fractions, indicating that the galaxy seems to be disrupted by their environment.
@@ 4.1 The role of environment in barred galaxies
This section will examine the effect of the environment on the bar structures of disc and lenticular descendants of barred galaxies. To accomplish this, we will first examine the merger histories of each sample. To increase our statistics in this section, we will analyse the lenticular and disc galaxies together. Fig 8 illustrates the fraction of descendants that belong to the barred disc and lenticular categories and have experienced major/minor mergers since redshift $z=1$. The study showcases the galaxy fractions with the lowest values of major and minor mergers (less than $0.28$ and $0.25$, respectively) for barred disc descendant galaxies that have maintained their bar structure. The maximum fraction of galaxies experiencing major mergers is observed when disc galaxies undergo bar dissolution, with an approximate value of $0.40$. The fraction of disc galaxies that have experienced a minor merger exhibits a consistent trend across all subsamples. In particular, galaxies that are descendants of barred discs and have retained their bar structure show the lowest fractions. Conversely, barred galaxies that lose their bar present the highest fractions. It should be noted that the fraction of galaxies experiencing a minor merger is relatively smaller in comparison to those undergoing major mergers. Our findings indicate a correlation between the dissolution of a bar and the merger histories of galaxies. Indeed, previous studies, have proposed that mergers can also lead to the weakening of a bar, especially if they have a contribution to the central part of the galaxy since can have the stronger gravitational potential of weakening the bar (e.g. Zana et al. 2018b; Ghosh et al. 2021).
@@
Regarding lenticular galaxies, there is an observed trend in contrast to what is typically observed in disc galaxies. In particular, galaxies that have retained their bars exhibit the highest fraction of galaxies that have undergone a significant merger ($>0.50$), while galaxies that have not developed bars demonstrate a comparable ratio ($0.40$). This increase may account for the observed rise in the mass percentage of the bulge and the disruption of the disc, as depicted in Fig 7, during later stages of barred lenticular descendants galaxies that have retained their bar structure. Because of the lower statistic of lenticular that lose their bar, we do not have information about the impact of minor merges. This finding implies that the environmental factors surrounding the dissolution of a bar after redshift $z=1$ may play a significant role.
@@
Figure 8: The fraction of barred galaxies that experienced a major/minor merger since $z=1$ for the disc and lenticular descendants. Errorbars represent the poison errors of each sample. The highest fraction of major/minor mergers corresponds to these subsamples where a bar structure dissolves, whereas the quiet histories are from disc galaxies that preserved their bar.
@@
Figure 9: Distributions of $\eta_{\rm dyn,db}$ defined in Eq 1 for barred galaxies that lose their bar and undergo a major merger. This expression asses if the disc galaxy experienced a recent merger or not at the moment of bar dissolution, and values between $-2,2$ indicate the galaxy is still influenced by a recent merger. 50 per cent of the disc galaxies that lose their bar have experienced a recent merger. Lenticular are not included because of the lower statistic number of galaxies that lose their bar.
@@
To closely examine the impact of major mergers in the galaxies that have retained their bars and compare them to those that have undergone bar dissolution, we calculate the temporal zone of merger influence, as stated in section 2.5. The quantification of this influence can be achieved through the parameter $\eta_{\rm dyn,db}$, as defined in Eq. 1. With this parameter, the discrepancy between the time of dissolution and the time of the most recent major/minor merger or the upcoming major/minor merger is computed relative to the dynamical time of the halo at the moment of bar dissolution. As a gentle reminder to the reader, it is important to note that negative values of $\eta_{\rm dyn,db}$ indicate the occurrence of a future merger, whereas positive values indicate that a merger has occurred in the past. The merger influence region during which the merger of the galaxy could potentially have an impact spans from negative two to positive two.
@@
To evaluate the impact of significant interactions on the disruption of a bar, we also incorporate the right panel of Fig 9. The presented figure illustrates the distribution of $\eta_{\rm dyn,db}$ values for galaxies that experience the loss of their bar and undergo a significant merger in close temporal proximity to the onset of bar strength weakening. As depicted in the diagram, it is observed that the barred disc galaxies undergoing a major merger experience a weakening of the bar, which finally leads to its disappearance. The suggested scenario is not applicable to lenticular galaxies, which we do not have an example. However, the weakening of the bar could be attributed to tidal interactions induced by flybys and potentially destroy them (e.g., Zana et al. 2018b).
@@
Figure 10: The median distance to a massive companion (mass ratio $=1/4$) of disc and lenticular descendant galaxies that lose their bar and do not undergo a major merger (green line) versus barred galaxies that did not undergo a major merger as a function of time since $z=1$. The shaded area corresponds to the $20^{\rm th}$ and $80^{\rm th}$ percentiles of the distribution. Galaxies that lose their bars are closer to their nearest massive companion than galaxies that retain their bars.
@@
Figure 11: The redshift evolution of the satellite fraction of barred discs and lenticular that preserved their bar and barred discs and lenticular that lose their bar. In general, the satellite fraction of lenticular and discs that lose their bar is higher in comparison to those that retain their bar.
@@
As a means of investigation, we compute the distance to the closest massive companions, as outlined in section 2.5, with the condition that the mass stellar ratio between the companion and the primary galaxy exceeds $1/4$. The calculation is performed for lenticular and disc galaxies that undergo bar dissolution and do not undergo a significant merger event, as a function of the lookback time. This is shown in Fig 10 where it is evident that the galaxies being examined display diverse interactions with massive companions at distances smaller than $100$ kpc, particularly during the early stages, near $z=1$. In order to establish a basis for comparison, we further compute the distance to the closest massive companion for barred disc and lenticular galaxies that have retained their bars and have not undergone significant merger events. We find that the distance to a massive galaxy is systematically higher in comparison with galaxies that lose their bar. Note that this occurs for lenticular and disc descendants, suggesting that tidal interactions are an important factor in the dissolution of bars. In fact, taking into account lenticular and disc galaxies, $50$ per cent of them become satellites before the bar weakens with a difference between the time of dissolution and the time of becoming a satellite being less than $1$ Gyr except for one case that is almost $2$ Gyrs, including the galaxies with mergers. This suggests that entry into a more dense environment contributes to the destruction of a bar. This is in agreement with a recent study of cluster galaxies from the extended WIde-field Nearby Galaxy-cluster Survey (OmegaWINGS) which found that galaxies in the outliers of the cluster undergoing tidal interactions due to massive companions destroying the bar (Tawfeek et al. 2022).
@@
We conclude this section by presenting the satellite fraction for disc and lenticular descendants in the last 8 Gyrs ($z=1$). Fig. 11 depicts the evolution of the satellite fraction for the barred discs and lenticulars that retain their bar and compares them to those that lose their bar. As is expected, disc galaxies that lose their bar, in general, live in denser environments and then present a higher satellite fraction ($0.50$) in comparison to barred disc and lenticular galaxies that retain their bar ($0.30$). This happens later in time in agreement that the dissolution of a bar could be triggered by tidal interactions by a close encounter or by ram pressure stripping due to torque forces in a dense environment. In conclusion, our results suggest galaxies that lose their bar are exposed to tidal interactions.

### 07 Feb 2024 v2

@@ 3.1 Merger histories and satellite fraction
In this subsection, we examine the merger histories of the three descendant samples (disc, lenticular, and spheroidal) in which mergers have been identified, as discussed in section 2.5.
We examine the cumulative distribution of major and minor mergers experienced by galaxies from redshift $z=1$ (corresponding to approximately 8 Gyr ago) to redshift $z=0$. The analysis is conducted on the descendant galaxy samples as depicted in Fig 4. Overall, descendant galaxies with a disc morphology have relatively quiet merger histories. Specifically, $85$ per cent of these galaxies have not undergone a major merger in the past $8$ Gyr. In contrast, only $38$ per cent and $36$ per cent of the descendant lenticular and spheroidal galaxies, respectively, are free of major mergers. This is in agreement with previous theoretical studies (e.g. see Fig.16 in Izquierdo-Villalba et al. 2019). Additionally, in the second panel of Fig. 4, we see that disc descendant galaxies that have undergone a significant merger, on average, experienced their last major merger event approximately $4.7$ Gyr ago, but lenticular and spheroidal descendant galaxies typically underwent their last large merger later on, at $3.7$ and $3.8$ Gyr ago, respectively.
Regarding the minor merger histories, no notable distinction is observed in the cumulative distribution across descendant galaxies of different types, with more than 30 per cent of them experiencing at least one minor merger. Specifically, approximately $62$ per cent of disc and lenticular galaxies and $66$ per cent of spheroidal descendant galaxies do not undergo minor mergers. The lenticulars underwent their most recent minor merger on average $3.9$ Gyr ago. In contrast, the disc and lenticular descendant galaxies exhibited earlier occurrences of minor mergers, thereby confirming the notion of relatively quiet evolutionary paths for these types of galaxies. The data presented in Table 2 illustrate the fraction of galaxies that have undergone at least one significant merger event since redshift $z=1$ across various samples. Additionally, the table provides the median values for the most recent major and minor merger events within each sample.
Figure 4: Merger histories of the three galaxy samples as specified in the legend. From left to right: Cumulative distribution of major mergers, cumulative distribution of the last major merger, cumulative distribution of the minor mergers, and cumulative distribution of the last minor merger. Overall, the descendant disc galaxies have more quiet major merger histories when compared to lenticular and spheroidal descendant galaxies.
Table 2: From left to right columns: Fraction of galaxies with at least one major merger since $z=1$, fraction of galaxies with at least one minor merger, and median lookback time of last major ($t_{\rm Mm}$[Gyr]) or minor ($t_{\rm mm}$[Gyr]) merger.
@@
One significant element that may be playing a role in this morphological transformation is the influence of tidal torques generated by the dense environment in which these galaxies are located (e.g. Martinez-Valpuesta et al. 2016; Łokas et al. 2016; Łokas 2021; Méndez-Abreu, Costantin, & Kruk 2023). This influence becomes particularly pronounced when these galaxies transition into denser environments. One way to gain insight into this phenomenon is by examining the evolution of the satellite fraction for each descendant sample, because satellite galaxies are susceptible to experiencing heightened levels of tidal torque, ram pressure stripping, and mergers.
@@
Figure 5 displays the fraction of satellites in the parent sample and each descendant sample. Overall, the satellite fraction increases with time. The highest increase is presented by the spheroidal descendants, with satellite galaxies accounting for $20$ per cent at redshift $z=1$ and expanding to $50$ per cent in the current epoch. Disc and lenticular descendants also present an increase, albeit more modest. Specifically, the fraction of galaxies classified as satellites is less than $10$ per cent at redshift $z=1$, and gradually increases to reach approximately 40 per cent. These galaxies are marked by a higher occurrence of interactions and a larger degree of exposure to tidal forces. Several processes in denser environments have the capacity to either disrupt or create a bar structure. This topic is further examined in the following sections.
@@
Figure 5: Redshift evolution of the satellite fraction for the parent sample, the disc sample, and the lenticular and spheroidal descendant samples. The satellite fraction increases with decreasing redshift, with the highest evolution presented for the spheroidal descendants.
4 Evolution of the $z=1$ barred galaxies
@@
In this section, we explore the evolution of the bar structures in the barred galaxies at $z=1$. For this, we focus on disc and lenticular galaxies. As discussed in the previous section, disc descendant galaxies, despite having relatively quiet merger histories in general, exhibit a diversity of merger histories and environments (see Fig 4), which must be taken into account when investigating the evolution of bar structures. For the lenticular galaxies, this diversity in the merger histories is higher. Each barred galaxy in the disc and lenticular descendants has been divided into two subsamples: (1) the descendants of barred galaxies at $z=1$ remain as barred galaxies at $z=0$; and (2) the descendants of barred galaxies at $z=1$ lose their bar, becoming unbarred galaxies at $z=0$.
@@
Figure 6: Evolution of barred disc galaxies divided into two subsamples: barred galaxies at $z=1$ that remain barred galaxies at $z=0$ (left column), and those that do not have their bar at $z=0$ (right column). Solid lines represent the median values and the shaded area corresponds to the $20^{\rm th}$ and $80^{\rm th}$ percentiles of the distribution.
@@
Figure 7: Evolution of lenticular descendant galaxies divided into two subsamples: Barred galaxies at $z=1$ that remain barred at $z=0$ (left column), and those that do not have their bar at $z=0$ (right column). The evolution of disc galaxies is shown in Fig. 6. Solid lines represent the median values and the shaded area corresponds to the $20^{\rm th}$ and $80^{\rm th}$ percentiles of the distribution. We note that here we only have one lenticular galaxy that loses its bar.
@@
In our analysis, which is summarised in Table 1, it is shown that a total of six galaxies, or $3$ per cent of the parent sample (the $z=1$ disc galaxies), lose the bar structure. On the other hand, one-third of the galaxies in our parent sample, namely $56$ barred galaxies or $30$ per cent (including bars only in lenticular and disc galaxies), maintain a bar structure at $z=0$. Upon examination of each descendant barred galaxy sample, a consistent pattern emerges: a minority of disc and lenticular descendent galaxies experience bar dissolution, with fractions of $4.7$ per cent and $1.1$ per cent, respectively.
4.1 Galaxy and bar properties across time
@@
The evolution of the properties of the bar structures is illustrated in Fig. 6 and Fig. 7 for the barred disc and lenticular galaxies, respectively. The first columns of the figures depict the evolution of the bar structures in the descendant galaxies that have retained their bars until the present epoch ($z=0$). We can appreciate that the bars that were already formed at $z=1$ formed at an early time (approximately $10$ Gyr ago), and the evolution of their bars is typical of that observed in isolated galaxies. Previous studies (Martinez-Valpuesta, Shlosman, & Heller 2006) showed that bars experience rapid growth during their dynamical phase, which lasts a few gigayears, after which their evolution becomes secular and remains relatively constant. However, there is a possibility that the bars may slightly weaken over time due to a buckling instability —as explained by Martinez-Valpuesta, Shlosman, & Heller (2006)—, and then remain constant in their secular evolution. This instability could potentially account for some of the fluctuations observed in the length of bars. Upon examining the morphological evolution depicted in the third and fourth rows of Fig. 6 and Fig. 7, it becomes evident that there is a lack of significant evolution in the thick-disc fraction. However, a marginal drop in the thin-disc fraction is observed (also seen in the evolution of the lenticular sample). This phenomenon is associated with the rise in the bulge mass fraction ($B/T$) of both classical bulges and pseudobulges during later times (see also Izquierdo-Villalba et al. 2022). This increase can likely be attributed to the presence of a bar, which facilitates the inflow of gas towards the central regions of galaxies (Spinoso et al. 2017; Donohoe-Keyes et al. 2019; George, Subramanian, & Paul 2019; Fraser-McKelvie et al. 2020) or may be the result of mergers; although in the following section we show that barred galaxies of the disc have the quietest merger histories. This increase in the $B/T$ is also found in lenticular galaxies, and the magnitude of this increase is greater. It is important to observe that the bars from the descendant barred disc galaxies that have retained their bar structure are more elongated than those in lenticular galaxies.
@@
The evolution of the bar structures that are dissolved is shown in the right columns of Fig. 6 and Fig. 7. These galaxies under consideration constitute a smaller fraction, specifically $4.7$ and $1.1$ per cent, of the total population of disc and lenticular descendants, respectively. We notice that these bars are young and small, and just recently formed before $z=1$. Regarding morphology, the disc mass fraction remains relatively stable at $z=0$ for the disc galaxy samples, with occasional changes likely attributed to either a major or a minor merger event. The bulge mass fraction exhibits a noticeable rise but with a significant degree of scatter. Additionally, it is important to note that the thick disc and pseudobulge components do not exhibit any signs of evolutionary change. In the case of the lenticular galaxies that lose their bar (only one case), not only is the bar structure disrupted but so is the thin disc, whereas there is no increase in the classical and pseudobulge mass fractions, indicating that the galaxy seems to be disrupted by its environment.
@@ 4.2 The role of environment in barred galaxies
In this section, we examine the effect of environment on the bar structures of disc and lenticular descendants of barred galaxies. To accomplish this, we first examine the merger histories of each sample. To increase our statistics in this section, we analyse the lenticular and disc galaxies together. Figure 8 illustrates the fraction of descendants that belong to the barred disc and lenticular categories and have experienced major or minor mergers since redshift $z=1$. Our study shows that the galaxy fractions with the lowest values of major and minor mergers (less than $0.28$ and $0.25$, respectively) are seen for barred disc descendant galaxies that have maintained their bar structure. The maximum fraction of galaxies experiencing major mergers is observed when disc galaxies undergo bar dissolution, with an approximate value of $0.40$. The fraction of disc galaxies that have experienced a minor merger exhibits a consistent trend across all subsamples. In particular, galaxies that are descendants of barred discs and have retained their bar structure show the lowest fractions. Conversely, barred galaxies that lose their bar present the highest fractions. It should be noted that the fraction of galaxies experiencing a minor merger is small in relation to those undergoing major mergers. Our findings indicate a correlation between the dissolution of a bar and the merger histories of galaxies. Indeed, previous studies proposed that mergers can also lead to the weakening of a bar, especially if they have a contribution to the central part of the galaxy, as the stronger gravitational potential can weaken the bar (e.g. Zana et al. 2018b; Ghosh et al. 2021).
@@
Regarding lenticular galaxies, there is an observed trend that is in contrast to what is typically observed in disc galaxies. In particular, galaxies that have retained their bars are the most likely to have undergone a significant merger ($>0.50$), while galaxies that have not developed bars demonstrate a comparable ratio ($0.40$). This increase may account for the observed rise in the mass percentage of the bulge and the disruption of the disc (as depicted in Fig 7) during later stages of the evolution of barred lenticular descendant galaxies that have retained their bar structure. Because we have a smaller sample size of lenticular galaxies that have lost their bar, we do not have information about the impact of minor merges. This finding implies that the environmental factors surrounding the dissolution of a bar after redshift $z=1$ may play a significant role.
@@
Figure 8: Fraction of barred galaxies that experienced a major or minor merger since $z=1$ for the disc and lenticular descendants. Error bars represent the poison errors of each sample. The highest fraction of major or minor mergers corresponds to the subsamples where a bar structure dissolves, whereas the quiet histories are from disc galaxies that preserved their bar.
@@
Figure 9: Distributions of $\eta_{\rm dyn,db}$ defined in Eq. 1 for barred galaxies that lose their bar and undergo a major merger. This expression assesses whether or not the disc galaxy experienced a recent merger at the moment of bar dissolution, and values between $-2$ and $2$ indicate that the galaxy is still influenced by a recent merger. Here, we see that 50 per cent of the disc galaxies that have lost their bar experienced a recent merger. Lenticular galaxies are not included because of the lower sample size we have for lenticular galaxies that have lost their bar.
@@
To closely examine the impact of major mergers in the galaxies that have retained their bars and compare them to those that have undergone bar dissolution, we calculated the temporal zone of merger influence, as stated in section 2.5. Quantification of this influence can be achieved through the parameter $\eta_{\rm dyn,db}$, as defined in Eq. 1. With this parameter, the discrepancy between the time of dissolution and the time of the most recent major or minor merger or the upcoming major or minor merger is computed relative to the dynamical time of the halo at the moment of bar dissolution. As a gentle reminder to the reader, it is important to note that negative values of $\eta_{\rm dyn,db}$ indicate the occurrence of a future merger, whereas positive values indicate that a merger has occurred in the past. The temporal range within which the merger may have an impact on the galaxy, relative to the dynamical time of the halo, spans from $[-2,2]$.
@@
To evaluate the impact of significant interactions on the disruption of a bar, we also incorporate the right panel of Fig 9, which illustrates the distribution of $\eta_{\rm dyn,db}$ values for galaxies that experience the loss of their bar and undergo a significant merger in close temporal proximity to the onset of bar weakening. As depicted in the diagram, it is observed that the barred disc galaxies undergoing a major merger experience a weakening of the bar, which finally leads to its disappearance. The suggested scenario is not applicable to lenticular galaxies, for which we do not have an example. However, the weakening of the bar could be attributed to tidal interactions induced by flybys and this could potentially destroy them (e.g. Zana et al. 2018b).
@@
Figure 10: Median distance to a massive companion (mass ratio $=1/4$) for disc and lenticular descendant galaxies that have lost their bar and have not undergone a major merger (green line) versus barred galaxies that have not undergone a major merger as a function of time since $z=1$. The shaded area corresponds to the $20^{\rm th}$ and $80^{\rm th}$ percentiles of the distribution. Galaxies that have lost their bars are closer to their nearest massive companion than galaxies that have retained their bars.
@@
Figure 11: Redshift evolution of the satellite fraction of barred discs and lenticular galaxies that retain their bar versus that of barred discs and lenticular galaxies that lose it. In general, the satellite fraction of lenticular and discs that lose their bar is higher in comparison to those that retain their bar.
@@
As a means of investigation, we computed the distance to the closest massive companions, as outlined in section 2.5, with the condition that the mass stellar ratio between the companion and the primary galaxy exceeds $1/4$. We performed this calculation for lenticular and disc galaxies that have undergone bar dissolution and have not experienced a significant merger event, and show the result as a function of lookback time. This is shown in Fig. 10, where it is evident that the galaxies being examined display diverse interactions with massive companions at distances of less than $100$ kpc, particularly during early stages, near $z=1$. In order to establish a basis for comparison, we further computed the distance to the closest massive companion for barred disc and lenticular galaxies that have retained their bars and have not experienced significant merger events. We find that the distance to a massive galaxy is systematically higher in comparison with galaxies that lose their bar. We note that this occurs for lenticular and disc descendants, suggesting that tidal interactions are an important factor in the dissolution of bars. More precisely, taking into account lenticular and disc galaxies, $50$ per cent of them become satellites before the bar weakens, with a difference between the time of dissolution and the time of becoming a satellite being less than $1$ Gyr, except for one case where this takes almost $2$ Gyr, including the galaxies with mergers. This suggests that entry into a more dense environment contributes to the destruction of a bar. This is in agreement with a recent study of cluster galaxies from the extended WIde-field Nearby Galaxy-cluster Survey (OmegaWINGS), where the authors found that galaxies situated on the periphery of a galaxy cluster are susceptible to tidal interactions, which have the potential to cause the destruction of the bar (Tawfeek et al. 2022).
@@
We conclude this section by presenting the satellite fraction for disc and lenticular descendants in the last 8 Gyr ($z=1$). Figure 11 depicts the evolution of the satellite fraction for the barred discs and lenticular galaxies that retain their bar and compares this to the evolution of the satellite fraction of those that lose their bar. As expected, disc galaxies that lose their bar, in general, live in denser environments and then present a higher satellite fraction ($0.50$) in comparison to barred disc and lenticular galaxies that retain their bar ($0.30$). This happens later in time, which is in agreement with the idea that the dissolution of a bar could be triggered by tidal interactions by a close encounter or by ram pressure stripping due to torque forces in a dense environment. In conclusion, our results suggest that galaxies that lose their bar are exposed to tidal interactions.